A basket contains 6 oranges and 4 tangerines. A sample of 3 is drawn. Find the probability that 2 are tangerines and one is an orange.
Answer by stanbon(48568)

A basket contains 6 oranges and 4 tangerines. A sample of 3 is drawn. Find the probability that 2 are tangerines and one is an orange.
----------
# of ways to get 2 tangerines= 4C2=6
# of ways to get one orange = 6C1=6
Therefore # of ways to get 2 tan and 1 or = 6*6=36
# of ways to pick 3 from the 10 fruit = 10C3=[10*9*8]/[1*2*3]=120
P(2Tan and 1 Orange)= 36/120=3/10
Cheers,
Stan H.